### v1

# AudioToolAgent: An Agentic Framework for Audio-Language Models

Abstract
Large Audio-Language Models (LALMs) perform well on audio understanding tasks but lack multi-step reasoning and tool-calling found in recent Large Language Models (LLMs). This paper presents AudioToolAgent, a framework that coordinates audio-language models as tools via a central LLM agent that accesses tool adapters for audio question answering and speech-to-text. The agent selects tools, asks follow-up questions, and compares outputs for verification. Experiments with MMAU, MMAR, and MMAU-Pro show state-of-the-art accuracy: up to 74.10% on MMAU, 68.80% on MMAR, and 57.96% on MMAU-Pro. Monte Carlo sampling for shapley values across 374 configurations identifies effective agent-tool combinations. The modular design allows integration of new tools and eliminates the use of data and training costs. Code and reproduction materials are available at: https://github.com/GLJS/AudioToolAgent.
Index Terms—  Audio-Language Models, Agentic Framework, Multi-Modal Audio Understanding, Reasoning, Tool-Calling
1 Introduction
Understanding and reasoning about audio is central to human cognition. Recent progress in transferring this capability to machines spans two areas: the advancement of Large Language Models (LLMs) with reasoning and tool-calling capabilities [1, 2, 3, 4, 5], and the development of Large Audio-Language Models (LALMs) for tasks such as audio captioning, audio question answering, and speech recognition [6, 7, 8, 9, 10, 11, 12, 13].
While recent LALMs perform well on audio benchmarks [14, 15], few can call multiple tools [10], whereas general LLMs excel at reasoning and using external tools but lack direct audio processing. This paper combines these strengths by enabling an LLM agent to use audio models as tools. This approach achieves more adaptive audio understanding than LALMs while maintaining the deeper reasoning capabilities of general LLMs.
Fig. 1: Example of a chatbot using the AudioToolAgent framework. The agent is a large language model that coordinates the tools; the tools are audio-language models.
This paper introduces AudioToolAgent, a framework that treats audio-language models as tools and uses a central agent to coordinate them. The agent, a text-only LLM, cannot process audio directly. Instead, the LLM receives the audio file path with a question or prompt and possible answers and delegates new instructions to LALMs (tools) to be able to understand the audio. The system prompt of the LLM contains instructions for the agent on how to use the tools. Because the framework reuses pretrained state-of-the-art models, the proposed framework needs no new datasets or training. Researchers can add both new public and local tools without architectural changes.
The agent receives an audio input, a question, and answer choices. The agent uses this information to reason about the task, upon which it calls tools to be able to answer the question or prompt. For speech, the agent prioritizes speech-to-text tools to transcribe the audio. For environmental sounds or music, the agent uses general audio models to gather information. AudioToolAgent asks follow-up questions, invokes tools iteratively, compares outputs, and verifies disagreements by continuing to call tools with different inputs to increase reliability. Figure 1 shows an example of the framework in a chatbot.
Fig. 2: Schematic overview of the AudioToolAgent framework. The framework has two components: a central agent and a set of tool models. The agent is a large language model that coordinates the tools; the tools are audio-language models.
The contributions of this work are twofold:
  1. 1.
    A modular architecture with an agent that coordinates audio-language tools through tool adapters. By utilizing pretrained foundational models without data or finetuning, this provides a cost-effective approach for state-of-the-art performance. Experiments on MMAU [14], MMAR [16] and MMAR-Pro [15] show that both closed-source and open-source versions of AudioToolAgent outperform prior models in several domains.
  2. 2.
    A benchmark that evaluates the effectiveness of different LLMs for reasoning on audio tasks and a benchmark that evaluates the effectiveness of different audio-language models as tools.
2 Related Work
Recent advances in large language models (LLMs) have resulted in agents that perform tool calling to solve tasks [17]. This began with GPT-3.5’s function calling [18] and includes the Model Context Protocol (MCP) [19], which standardizes interactions with external tools. This work uses the ReAct framework [20]. In ReAct, the agent first reasons about the task and then performs actions to solve it. The agent selects appropriate tools to answer the question. After receiving tool responses, the agent decides whether to make additional tool calls or answer the question with the information already gathered.
Recent developments include large multimodal models that integrate audio processing and agentic capabilities within a single architecture. Models such as Gemini 2.5 [5] and GPT-4o [21] handle both audio understanding and speech recognition while performing tool calling. These models can generate audio output, enabling real-time, end-to-end speech-to-speech interactions with tool use. Training these models costs substantial resources, and they remain closed-source, accessible only through API endpoints.
Another relevant work, StepAudio 2 [10], received explicit training for tool calling and benchmarking on four specific tools: audio search with multimodal RAG, date and time retrieval, weather search, and web search. This model processes audio input, performs tool calls, understands audio content, and generates speech output. Training this integrated model consumed 1.356 trillion tokens over 21 days [10]. In contrast, the AudioToolAgent framework eliminates this training cost by coordinating existing pre-trained models.
Similarly, other works with the same name include AudioAgent [22], which uses audio attributes to optimize prompts via a fine-tuned LLM for audio tools, and the Audio-Agent framework [23], which uses an LLM to orchestrate audio generation and editing. The current work differs by using a text-only agent that delegates audio understanding to specialized tools without fine-tuning. Instead of using fixed classifiers, AudioToolAgent queries multiple interchangeable tools and cross-checks their outputs for verification.
3 Methodology and Experimental Setup
3.1 Framework Overview
The agent, a reasoning model, receives an audio file path and task description and selects tools to produce the output. It accesses audio signals only through tools. The agent identifies suitable tool calls for the task, then invokes them through structured tags: <tool_call> to initiate a request and </tool_call> to conclude it. Within these tags, the agent specifies the target tool, audio file path, and prompt. Each tool’s output enters the agent’s context, enabling it to reason and invoke additional tools as needed. The tool set includes audio understanding and speech recognition tools. Figure 2 shows a schematic visualization. To prevent runaway loops, each agent can invoke a maximum of 20 tool calls. In AudioToolAgent, the agent typically makes 5-10 calls, depending on the configuration.
Instructions in the system prompt help the agent to issue follow-up tool calls when outputs conflict or remain ambiguous, and to gather targeted evidence rather than guessing. The prompt also directs the agent to resolve conflicting or ambiguous tool outputs through targeted follow-up queries and to format answers within <answer> tags for parsing. The system prompt starts as follows:
System Prompt You are an expert audio analyst with access to specialized tools. Answer the question given. Put the answer between <answer> and </answer> tags. If the question is multiple choice, there is always just one choice correct. If the tool says it can’t listen to audio, try invoking the tool again. Use as many different tools as needed to answer the question, even using the same tool multiple times if needed. If initial tool outputs are conflicting or ambiguous, do not guess; instead, you must generate specific, follow-up tool calls to isolate the point of disagreement and gather more detailed evidence. The following tools are available …
In this framework, all tools connect via HTTP API tool adapters for modularity. This includes public endpoints for proprietary models and self-hosted endpoints for open-source models, running on either vLLM [24] or Transformers [25]. To demonstrate the framework’s versatility, the implementation offers two configurations:
  • AudioToolAgent: This configuration uses a proprietary agent and closed-source tools, accessed through public API endpoints to maximize performance. The agent is the GPT5 model [4], and the tool suite includes GPT-4o [21], Gemini 2.5 Flash [5], Voxtral [12], Qwen2.5 Omni [6], and Audio Flamingo 3 [11].
  • AudioToolAgent-Open: As the primary model, this configuration uses an open-source agent with high-performing open-source audio tools to balance performance with self-hosting capabilities. The agent is DeepSeek V3.1 [1], and the tool suite includes Whisper [13], Voxtral [12], Qwen2.5 Omni [6], Audio Flamingo 3 [11], and DeSTA 2.5 [9].
3.2 Evaluation Setup
The study evaluated AudioToolAgent and AudioToolAgent-Open on three benchmarks: Massive Multi-Task Audio Understanding (MMAU) [14], MMAR [16] and MMAR-Pro [15]. The MMAU benchmark includes 10,000 audio clips for multi-task audio understanding and reasoning, with 1,000 in the test-mini split and 9,000 in the test split. The experiments used only the test-mini split to reduce costs. The MMAR benchmark tests deep reasoning capabilities with 1,000 audio-question-answer triplets requiring multi-step reasoning across modalities. The MMAU-Pro benchmark measures audio intelligence using 5,304 instances (one audio example was broken) containing human expert-generated question-answer pairs across speech, sound, music, and combinations.
4 Results
Dataset Models Results
MMAU test-mini [14] Sound $|$ Music $|$ Speech $|$ Average Closed Source
GPT-4o Audio [21] 64.56 $|$ 56.29 $|$ 66.67 $|$ 62.50
Gemini 2.5 Pro [5] 75.08 $|$ 68.26 $|$ 71.47 $|$ 71.60
Omni-R1 [26] 81.70 $|$ 73.40 $|$ 76.00 $|$ 77.00
Step-Audio 2 [10] 83.48 $|$ 73.65 $|$ 76.88 $|$ 78.00
AudioToolAgent 73.57 $|$ 69.16 $|$ 79.57 $|$ 74.10
Open Source
Audio Reasoner [27] 67.87 $|$ 69.16 $|$ 66.07 $|$ 67.70
Kimi-Audio [8] 75.68 $|$ 66.77 $|$ 62.16 $|$ 68.20
Qwen2.5-Omni [6] 78.10 $|$ 65.90 $|$ 70.60 $|$ 71.50
Step-Audio 2 mini [10] 76.28 $|$ 71.56 $|$ 71.47 $|$ 73.20
Audio Flamingo 3 [11] 79.58 $|$ 66.77 $|$ 66.37 $|$ 73.30
AudioToolAgent-Open 78.08 $|$ 69.67 $|$ 75.08 $|$ 74.20
MMAR [16] Sound $|$ Music $|$ Speech $|$ Sound-Music $|$ Sound-Speech $|$ Music-Speech $|$ Sound-Music-Speech $|$ Average Closed Source
GPT-4o Audio [21] 53.94 $|$ 50.97 $|$ 70.41 $|$ 63.64 $|$ 72.48 $|$ 62.20 $|$ 75.00 $|$ 63.50
Gemini 2.0 Flash [28] 61.21 $|$ 50.97 $|$ 72.11 $|$ 81.82 $|$ 72.48 $|$ 65.85 $|$ 70.83 $|$ 65.60
Omni-R1 [26] 67.30 $|$ 51.50 $|$ 64.30 $|$ 45.50 $|$ 70.20 $|$ 64.60 $|$ 70.80 $|$ 63.40
AudioToolAgent 61.81 $|$ 51.94 $|$ 77.55 $|$ 72.72 $|$ 76.61 $|$ 71.96 $|$ 70.83 $|$ 68.80
Open Source
Audio Reasoner [27] 43.64 $|$ 33.50 $|$ 32.99 $|$ 45.45 $|$ 42.66 $|$ 31.71 $|$ 25.00 $|$ 36.80
Qwen2.5-Omni [6] 58.79 $|$ 40.78 $|$ 59.86 $|$ 54.55 $|$ 61.93 $|$ 67.07 $|$ 58.33 $|$ 56.70
AudioToolAgent-Open 59.39 $|$ 45.63 $|$ 67.34 $|$ 54.55 $|$ 70.64 $|$ 59.76 $|$ 70.83 $|$ 61.70
MMAU-Pro [15] Sound $|$ Music $|$ Speech $|$ Sound-Music $|$ Speech-Music $|$ Speech-Sound $|$ Sound-Music-Speech $|$ Spatial $|$ Voice $|$ Multi-Audio $|$ Open-ended $|$ Instruction-Following $|$ Average Closed Source
GPT4o Audio [21] 44.70 $|$ 63.10 $|$ 68.20 $|$ 40.40 $|$ 43.50 $|$ 62.50 $|$ 57.10 $|$ 21.40 $|$ 57.50 $|$ 32.60 $|$ 43.20 $|$ 82.50 $|$ 52.50
Gemini-2.5 Flash [5] 51.90 $|$ 64.90 $|$ 73.40 $|$ 42.80 $|$ 58.70 $|$ 61.30 $|$ 42.80 $|$ 36.30 $|$ 71.70 $|$ 21.20 $|$ 67.50 $|$ 95.10 $|$ 59.20
AudioToolAgent 33.14 $|$ 63.47 $|$ 73.74 $|$ 26.00 $|$ 50.00 $|$ 54.55 $|$ 57.14 $|$ 30.15 $|$ 70.69 $|$ 57.21 $|$ 73.31 $|$ 86.21 $|$ 57.96
Open Source
Audio-Reasoner [27] 34.20 $|$ 50.10 $|$ 44.00 $|$ 26.00 $|$ 36.90 $|$ 43.20 $|$ 28.60 $|$ 20.30 $|$ 43.40 $|$ 22.60 $|$ 38.60 $|$ 43.40 $|$ 39.50
Kimi-Audio [8] 46.00 $|$ 57.60 $|$ 52.20 $|$ 46.00 $|$ 54.30 $|$ 48.90 $|$ 42.80 $|$ 43.70 $|$ 50.60 $|$ 17.20 $|$ 34.50 $|$ 42.30 $|$ 46.60
Audio Flamingo 3 [11] 55.90 $|$ 61.70 $|$ 58.80 $|$ 40.00 $|$ 41.30 $|$ 47.70 $|$ 57.10 $|$ 26.80 $|$ 58.60 $|$ 26.00 $|$ 44.20 $|$ 33.30 $|$ 51.70
Qwen2.5-Omni [6] 47.60 $|$ 61.50 $|$ 57.40 $|$ 40.00 $|$ 53.20 $|$ 60.20 $|$ 28.50 $|$ 41.20 $|$ 60.00 $|$ 24.30 $|$ 52.30 $|$ 61.30 $|$ 52.20
AudioToolAgent-Open 42.79 $|$ 64.39 $|$ 67.90 $|$ 36.00 $|$ 45.65 $|$ 54.55 $|$ 57.14 $|$ 33.54 $|$ 61.72 $|$ 23.49 $|$ 72.17 $|$ 64.37 $|$ 55.68
Table 1: Comparison of AudioToolAgent with baseline models on MMAU, MMAR, and MMAU-Pro benchmarks. AudioToolAgent achieves state-of-the-art performance across all evaluation metrics. All baseline model scores are copied from their respective original works and/or benchmark evaluations. Self-proposed, no code or API available to verify.
Table 1 summarizes the results. AudioToolAgent-Open outperforms all open-source models on average across the three benchmarks, including the individual tools it uses. AudioToolAgent outperforms most closed-source models on average, even when some models could not be used as the tools of AudioToolAgent due to no API availability.
Performance gains appear most pronounced in the Speech portions of the benchmarks (see Speech columns in Table 1). The automatic speech recognition tools in both AudioToolAgent and AudioToolAgent-Open explain this improvement. While other approaches use a single multimodal model trained for both speech recognition and audio understanding, AudioToolAgent invokes ASR models like Whisper [13] and Voxtral [12] for accurate transcription.
4.1 Ablation Study
To identify the most effective configuration for AudioToolAgent, an ablation study examined 10% of the MMAU test-mini split (100 examples), analyzing agents and tools separately.
4.1.1 Agents
To evaluate the LLMs capable of tool calling, the experiments used a fixed set of tools - the same tools from the open-source AudioToolAgent configuration (see Section 3.1). Each evaluation ran five independent tests per agent with different random seeds and reported the mean. This approach accounts for accuracy variations from non-deterministic inference even with fixed seeds, partly due to vendor-recommended decoding defaults such as nonzero temperature. Figure 3 shows the tested tools on the y-axis.
Figure 3 visualizes the agent ablation. Inspired by Omni-R1 [26], which showed that text-only models perform well on audio reasoning tasks, the black vertical tick on the bar plot shows each LLM’s performance without audio capabilities, which still scores well on the benchmark. The dots represent individual evaluations, with the horizontal colored bar showing the average across 5 runs.
0.350.350.40.40.450.450.50.50.550.550.60.60.650.650.70.70.750.750.80.80.850.85GLM 4.5 [2]Hermes 4 405B [29]Gemini 2.5 Flash [5]Qwen 3 Next 80B A3B [3]Qwen 3 235B A22B [3]Qwen 3 Coder 480B A35B [30]GPT-5 Mini [4]GPT-5 [4]Claude Sonnet 4.0 [31]Kimi K2 [32]Deepseek V3.1 [1]Accuracy
Fig. 3: Per-model accuracies (dots; 5 runs) and mean accuracy (horizontal bar) on a subset of the MMAU test-mini split. Vertical black ticks denote the corresponding ALM without any tools.
Deepseek V3.1 outperforms all other LLMs with a mean accuracy of 0.784, followed by Kimi K2 (0.766), Claude Sonnet 4 (0.762) and GPT-5 (0.748). Based on these results, AudioToolAgent-Open uses Deepseek V3.1 and AudioToolAgent uses GPT-5. GPT-5 was chosen over Claude Sonnet 4 because it costs less in the configuration: low settings for reasoning effort and verbosity were maintained to reduce costs.
4.1.2 Tools
To quantify each tool’s contribution to system performance, this work estimates Shapley values using a two-stage, Monte Carlo approximation. In the first stage, the method generates multiple sampled permutations of the available tools. For each permutation, the system evaluates performance on the 100 examples from the MMAU subset. In the second stage, the approach calculates the final Shapley values by considering only combinations of two or more tools to determine each tool’s marginal contribution. The final Shapley value for each tool represents the average of these marginal contributions across all sampled permutations. Figure 4 displays these values. The analysis clusters the tested tools into 4 categories, shown on the y-axis of Figure 4.
-0.050.000.050.100.15Tavily Web Search [33]DuckDuckGo Web Search [34]StepAudio 2 [10]Parakeet TDT v3 [35]Granite Speech [36]Speaker Diarization [37]Phi-4 Multimodal [38]Gemma 3n [39]GPT4o-mini [21]DeSTA2.5 [9]Voxtral [12]Whisper v3 Turbo [13]Kimi Audio [8]AudSemThinker [40]Qwen2Audio [7]Gemini 2.5 Flash [5]AudioFlamingo 3 [11]Qwen2.5 Omni [6]Shapley valueAudio UnderstandingSpeech-to-TextSpeaker DiarizationAuxiliary
Fig. 4: Shapley values per feature with standard error bars computed over 374 runs. Bars indicate the contribution of the audio-language model or API as tool to the overall system performance. Error bars are black.
Qwen2.5 Omni [6] provides the highest contribution, followed by Audio Flamingo 3 [11] and Gemini 2.5 Flash [5]. The top-performing tools were incorporated in the configurations of this work (see Section 3.1) with two exceptions: Qwen2Audio [7] processes only 30 seconds of audio and audios in benchmarks are often longer, and AudSemThinker [40] shares its architecture with Qwen2.5 Omni.
4.2 Discussion and Future Work
AudioToolAgent’s performance depends on its underlying audio tools, creating both opportunities and challenges for future development. When tools produce inaccurate outputs, the agent may propagate these errors. To mitigate this, the agent uses a cross-validation approach, verifying information across multiple tools. By comparing answers to direct questions, the agent reduces reliance on single tool responses. Future work should explore advanced consensus mechanisms and uncertainty quantification to improve robustness against tool errors.
A practical limitation of AudioToolAgent is speed. Using multiple separate tools creates longer processing times than a single audio model. The agent calls tools sequentially and waits for each result, adding overhead. Running AudioToolAgent on separate machines mitigates this issue. Future research should focus on training agents to select optimal tool subsets for specific tasks, improving efficiency.
We tested the framework with web search integration using both DuckDuckGo API and Tavily’s proprietary search API (see Figure 4). Ablation studies revealed no consistent improvements, likely because the benchmarks focus on audio content and general information. These benchmarks mostly require historical information and common knowledge facts. Nevertheless, web search integration remains promising for real-world applications where external knowledge retrieval enhances performance. The tests also included the ability to extract parts of the audio and use them as input to the tools, but this did not improve performance. Future work should explore this approach. Expanding the tool ecosystem to include audio retrieval, audio generation, and analysis tools offers another valuable research direction.
Our evaluation approach uses existing benchmark results from model authors and benchmark maintainers (see Table 1) rather than reproducing all baseline numbers independently. This decision was made based on practical factors: high inference costs, API availability limitations, and the need for model-specific optimization to achieve peak performance. The unverified entries are marked where applicable.
5 Conclusion
This paper introduced AudioToolAgent, a framework for multimodal audio understanding and reasoning where a central agent coordinates audio-language models as tools. The combination of GPT-4o, Gemini 2.5 Flash, Voxtral, Qwen2.5 Omni and Audio Flamingo 3 orchestrated by GPT-5 outperforms prior models on the MMAU, MMAR and MMAR-Pro benchmarks.
This framework establishes a new paradigm that combines the strengths of ALMs and LLMs. The ablation studies identified Qwen2.5 Omni and AudioFlamingo 3 as the most effective audio tools. Among LLMs, DeepSeek V3.1 and Kimi K2 demonstrated superior performance as orchestrating agents. This hybrid approach combines the audio processing of ALMs with the reasoning strengths of LLMs, creating a more flexible and powerful system than either model type could achieve alone.
This work opens several promising research directions. Future work includes expanding to other tasks and developing learned tool selection policies, improving the speed of the framework, and exploring the use of web search and other analysis tools.
6 Acknowledgment
This work was supported by the Dutch Research Council (NWO 406.20.GO.030 to Prof. Elia Formisano), the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-14224, Data Science Research Infrastructure (DSRI; Maastricht University) and the Dutch Province of Limburg.